A 50 kg crate rests on a level floor. Solve for the coefficient of static friction, if the static friction is 298.9 N.
1
Expert's answer
2020-11-17T11:30:03-0500

We can find the coefficient of static friction from the formula:


"F_{s.fr.}=\mu_sN=\mu_smg,""\mu_s=\dfrac{F_{s.fr.}}{mg}=\dfrac{298.9\ N}{50\ kg\cdot 9.8\ \dfrac{m}{s^2}}=0.61"

Answer:

"\mu_s=0.61"